# Attribute Inference Attacks for Federated Regression Tasks

### v1

@@ 1 Introduction
Figure 1: Average performance of different AIAs when four clients train a neural network through FedAvg with 1 local epoch and batch size 32. Each client stores $|{\mathcal{D}}_{c}|$ data points randomly sampled from ACS Income dataset (Ding et al. 2024). The adversary infers the gender attribute of every data sample held by the client given access to the released (public) information.
@@
  • We provide an analytical lower bound for model-based AIA accuracy in the least squares regression problem. This result motivates the adversary’s strategy to approximate the client’s optimal local model in federated regression tasks. (Sec. 3).
@@ 2.1 Federated learning
2:     $s$ selects a subset of the clients $\mathcal{C}_{s}^{t}\subseteq{\mathcal{C}}$,
3:     $s$ broadcasts the global model $\theta^{t}$ to $\mathcal{C}_{s}^{t}$,
4:     $s$ waits for the updated models $\theta_{c}^{t}$ from every client $c\in\mathcal{C}_{s}^{t}$,
@@
2:  for each local epoch $e$ from 1 to $E$ do
3:     for batch $b\in\mathcal{B}$ do
5:        $k\leftarrow k+1$
6:  Return $\theta_{c}^{t}(K)$, where $K=E\lceil\frac{S_{c}}{B}\rceil$
Let $\theta^{*}=\arg\min_{\theta\in\mathbb{R}^{d}}{\mathcal{L}}(\theta)$ be a global optimal model, i.e., a minimizer of Problem (2.1). A general framework to learn such a global model in a federated way is shown in Algorithm 1; it generalizes a large number of FL algorithms, including FedAvg (McMahan et al. 2017), FedProx (Li et al. 2020), and FL with different client sampling techniques (Nishio and Yonetani 2019; Chen, Horvath, and Richtarik 2020; Jee Cho, Wang, and Joshi 2022). The model $\theta^{T}$—the output of Algorithm 1—is the tentative solution of Problem (2.1). Its performance depends on the specific FL algorithm, which precises how clients are selected in line 2, how the updated local models are aggregated in line 5, and how the local update rule works in line 8. For example, in FedAvg, clients are selected uniformly at random among the available clients, the local models are averaged with constant weights, and the clients perform locally multiple stochastic gradient steps (Algorithm 2).
@@ 2.2 Threat Model
Honest-but-curious adversary
We describe first an honest-but-curious adversary, [^1] In what follows, we refer to the client using female pronouns and the adversary using male pronouns, respectively. which is a standard threat model in existing literature (Melis et al. 2019; Geiping et al. 2020; Yin et al. 2021; Nasr, Shokri, and Houmansadr 2019), including the FL one (Kairouz et al. 2021, Table 7), (Lyu and Chen 2021; Chen et al. 2022). This passive adversary, who could be the server itself, is knowledgeable about the trained model structure, the loss function, and the training algorithm, and may eavesdrop on communication between the attacked client and the server but does not interfere with the training process. For instance, during training round $t$, the adversary can inspect the messages exchanged between the server and the attacked client (denoted by $c$), allowing him to recover the parameters of the global model $\theta^{t}$ and the updated client model $\theta_{c}^{t}(K)$ (Algorithm 2). Let ${\mathcal{T}_{c}}\subseteq\{t|c\in\mathcal{C}_{s}^{t},\forall t\in{0,\dots,T-1}\}$ denote the set of communication rounds during which the adversary inspects messages exchanged between the server and the attacked client and ${\mathcal{M}}_{c}=\{(\theta^{t},\theta_{c}^{t}),\forall t\in{\mathcal{T}_{c}}\}$ denote the corresponding set of messages.
@@
Malicious adversary
We also consider a stronger active adversary who can interfere with the training process. Specifically, this adversary can modify the messages sent to the clients and have the clients update models $\theta^{t}$ concocted to reveal more private information. Let $\mathcal{T}^{a}_{c}$ be the set of rounds during which the adversary attacks client $c$ by sending malicious model $\theta^{t}$. As above, the adversary could be the server itself. This adversary has been widely considered in the literature on reconstruction attacks (Wen et al. 2022; Boenisch et al. 2023) and membership inference attacks (Nguyen et al. 2023; Nasr, Shokri, and Houmansadr 2019). Some studies have also explored the possibility of a malicious adversary modifying the model architecture during training (Fowl et al. 2022; Zhao et al. 2023), even though such attacks appear to be easily detectable. In this paper, we do not allow the adversary to modify the model architecture.
For simplicity, we will refer to these two adversaries as passive and active adversaries, respectively, throughout the rest of the paper.
@@ 2.3 Attribute inference attack (AIA) for FL
AIA leverages public information to deduce private or sensitive attributes (Kasiviswanathan, Rudelson, and Smith 2013; Lyu and Chen 2021; Chen et al. 2022; Fredrikson et al. 2014; Yeom et al. 2018). For example, an AIA could reconstruct a user’s gender from a recommender model by having access to the user’s provided ratings. Formally, each input $\mathbf{x}_{c}(i)$ of client $c$ consists of public attributes $\mathbf{x}_{c}^{p}(i)$ and of a sensitive attribute $s_{c}(i)$. The target value, assumed to be public, is denoted by $y_{c}^{p}(i)$. The adversary, having access to $\{(\mathbf{x}_{c}^{p}(i),y_{c}^{p}(i)),i=1,...,S_{c}\}$ and ${\mathcal{M}}_{c}$, aims to recover the sensitive attributes ${s_{c}(i)}$. [^2] In  (Fredrikson et al. 2014; Yeom et al. 2018), the adversary possesses additional information, including estimates of the marginals or the joint distribution of the data samples. However, in this paper, we do not consider such a more powerful adversary.
Existing gradient-based AIA for FL
Chen et al. (2022) present AIAs specifically designed for the FL context and both passive and active adversaries. The central idea involves identifying sensitive attribute values that yield virtual gradients closely resembling the client’s model updates—referred to as pseudo-gradients—in terms of cosine similarity. Formally, the adversary solves the following optimization problem:
@@
Chen et al. (2022) assume that the sensitive attributes are categorical and discrete random variables. Nevertheless, problem (2) can be solved efficiently using a gradient method with the reparameterization trick and the Gumbel softmax distribution (Jang, Gu, and Poole 2017). From (2), we observe that, since gradients incorporate information from all samples, the attack performance deteriorates in the presence of a large local dataset. For example, the attack accuracy almost halves on the Genome dataset for the classification task when the client’s local dataset size increases from 50 to 1000 samples (Lyu and Chen 2021, Table 8). Our experiment (Figure 1) on a regression task corroborates this finding: when the local dataset size increases from 64 to 256, the attack accuracy drops from $60\%$ to the level of random guessing.
Model-based AIA
@@
Below we provide theoretical guarantees (Prop. 1) for the accuracy of the model-based AIAs to least squares regression problems. Our theoretical result supports that in an FL setting, the adversary may advantageously start by estimating the client’s optimal local model, i.e., the model that minimizes the client’s empirical loss.
3 Model-based AIA guarantees for least squares regression.
@@
It is easy to check that the problem can be equivalently solved through the following two-step procedure. First, the adversary computes the vector of real values:
@@ 4.1 Passive approach for linear least squares
We consider that clients cooperatively train a linear regression model with quadratic loss. We refer to this setting as a federated least squares regression. The attack is detailed in Alg. 3 and involves a single matrix computation (line 3) after the exchanged messages ${\mathcal{M}}_{c}$ have been collected. $\mathbf{A}^{\dagger}$ represents the pseudo-inverse of matrix $\mathbf{A}$. Theorem 1 provides theoretical guarantees for the distance between the reconstructed model and the optimal local one, when the model is trained through FedAvg (McMahan et al. 2017) with batch size $B$ and local epochs $E$. The formal statement of the theorem and its proof are in Appendix A.1.
@@
Consider a federated least squares regression with a large number of clients and assume that i) client $c$ has $d$-rank design matrix $\mathbf{x}_{c}\in\mathbb{R}^{S_{c}\times d}$, ii) she updates the global model through stochastic gradient steps with sub-Gaussian noise with scale $\sigma$ and small learning rate $\eta$, and iii) the local data distribution of the targeted client is different from the global data distribution. By eavesdropping on $n_{c}>d$ message exchanges between client $c$ and the server, the error of the reconstructed model $\hat{\theta}^{*}_{c}$ of Algorithm 3 is upper bounded w.p. $\geq 1-\delta$ when $\eta\leq\frac{S_{c}}{2\lambda_{\max}(\mathbf{x}_{c}^{T}\mathbf{x}_{c})}$ and
$\left\lVert\hat{\theta}^{*}_{c}-\theta_{c}^{*}\right\rVert_{2}=\mathcal{O}\left(\eta\sigma d\sqrt{E\left\lceil\frac{S_{c}}{B}\right\rceil\frac{d+1+\ln\frac{2d}{\delta}}{n_{c}\cdot\underline{\lambda}}}\right).$ (6)
When the batch size is equal to the local dataset size, Alg. 3 exactly reconstructs the optimal local model (proof in Appendix A.2):
@@
Consider a federated least squares regression and assume that client $c$ has $d$-rank design matrix and updates the global model through full-batch gradient updates (i.e., $B=S_{c}$) and an arbitrary number of local epochs. Once the adversary eavesdrops on $d+1$ communication exchanges between client $c$ and the server, it can recover the client’s optimal local model.
@@
Toy example illustration
Here, we illustrate the performance of our Alg. 3 on a toy dataset detailed in Appendix B.2. Figure 2 (left) demonstrates that as the batch size increases, the reconstructed model is closer to the optimal local model (as shown in our Theorem 1). Moreover, since the model overfits more the dataset (with smaller loss), the AIA accuracy increases (Figure 2 (right)).
@@ 4.2 Active approach
The details of this attack are outlined in Alg. 4. We observe that the adversary does not need to systematically modify all messages sent to the target client $c$ but can modify just a handful of messages that are not necessarily consecutive. This contributes to the difficulty of detecting the attack.
@@ 5.2 FL training and attack setup
In all the experiments, each client follows FedAvg (Alg. 2) to train a neural network model with a single hidden layer of 128 neurons, using ReLU as activation function. The number of communication rounds is fixed to $T=\lceil 100/E\rceil$ where $E$ is the number of local epochs. Each client participates to all rounds, i.e., $\mathcal{T}_{c}=\{0,...,T-1\}$. The learning rate is tuned for each training scenario (different datasets and number of local epochs), with values provided in Appendix B.4. The passive adversary may eavesdrop all the exchanged messages until the end of the training. The active adversary launches the attack after $T$ rounds [^4] We also examine the impact of the attack’s starting round. The results are presented in the Appendix C.2. for additional $\lceil 10/E\rceil$ and $\lceil 50/E\rceil$ rounds. Every attack is evaluated over FL trainings from 3 different random seeds. For Medical dataset, the adversary infers whether a person smokes or not. For Income-L and Income-A datasets, the adversary infers the gender. The AIA accuracy is the fraction of the correct inferences over all the samples. We have also conducted experiments for federated least squares regression on the same datasets. The results can be found in Appendix C.1.
@@ 5.3 Baselines and our attack implementation
Gradient-based:
We compare our method with the (gradient-based) SOTA (Sec. 2.3). The baseline performance is affected by the set of inspected rounds $\mathcal{T}$ considered in (2). We select the inspected rounds $\mathcal{T}$ based on two criteria: the highest cosine similarity and the best AIA accuracy. In a real attack, the adversary is not expected to know the attack accuracy beforehand. Therefore, we refer to the attack based on the highest cosine similarity as Grad and to the other as Grad-w-O (Gradient with Oracle), as it assumes the existence of an oracle that provides the attack accuracy. The details for the tuning of $\mathcal{T}$ and other hyper-parameter settings can be found in Appendix  B.4.
Our Attacks:
Our attacks consist of two steps: 1) reconstructing the optimal local model, and 2) executing the model-based AIA (3). For the first step, a passive adversary uses the last-returned model from the targeted client, while an active adversary executes Alg. 4. The details of the hyperparameter settings can be found in Appendix B.4 .
Model-based with Oracle (Model-w-O):
@@ 5.4 Experimental results
Table 1: The AIA accuracy over all clients’ local datasets evaluated under both honest-but-curious (passive) and malicious (active) adversaries across Income-L, Income-A, and Medical FL datasets (Sec. 5.1). The standard deviation is evaluated over three FL training runs with different random seeds. All clients run FedAvg with 1 epoch and batch size 32. For Income-L, we consider the dataset with $40\%$ of data heterogeneity (Appendix B.3).
From Table 1, we see that our attacks outperform gradient-based ones in both passive and active scenarios across all three datasets. Notably, our passive attack achieves improvements of over $15$ and $8$ percentage points (p.p.) for the Income-L and Medical datasets, respectively. Even when the gradient-based method has access to an oracle, our passive attacks still achieves higher accuracy on two datasets and comes very close on Income-A. When shifting to active attacks, the gains are even more substantial. For instance, when the attack is active for 50 rounds, we achieve gains of 13, 22, and 5 percentage points (p.p.) in Income-L, Income-A, and Medical, respectively, over Grad-w-O, and even larger gains over the more realistic Grad. Furthermore, the attack accuracy reaches the performance expected from an adversary who knows the optimal local model. Interestingly, while our attacks consistently improve as the adversary’s capacity increases (moving from a passive attacker to an active one and increasing the number of rounds of the active attack), this is not the case for gradient-based methods.
@@
Figure 3 (right) shows that under passive attacks, our approach is not sensitive to the number of local epochs, whereas Grad’s performance deteriorates to the level of random guessing as the number of local epochs increases to 5. Interestingly, active Grad significantly improves upon passive Grad as the number of local epochs increases. While our active approach progressively performs slightly worse, it still maintains an advantage of over 10 p.p. compared to Grad.
@@
The figure on the right shows how each client’s dataset size impacts individual (active) AIA accuracy in Income-A dataset. The experiment setting is the same as in Table 1 with 50 active rounds. [Uncaptioned image] We observe that for all methods, clients with smaller datasets are more vulnerable to the attacks, because the models overfit the dataset more easily. Moreover, for clients with more than 20000 data points, Grad provides an attack performance close to random guessing.
@@
To mitigate privacy leakage, we apply a federated version of DP-SGD (Abadi et al. 2016), which provides sample-level differential privacy guarantees (Dwork et al. 2006; Choudhury et al. 2019). Our experiments show that even with this defense mechanism in place, our approach still outperforms the baselines in most of the scenarios. The results are in Appendix C.4.

### v2

@@ 1 Introduction
Figure 1: Average performance of different AIAs when four clients train a neural network through FedAvg with 1 local epoch and batch size 32. Each client stores $S_{c}$ data points randomly sampled from ACS Income dataset (Ding et al. 2024). The adversary infers the gender attribute of every data sample held by the client given access to the released (public) information.
@@
  • We provide an analytical lower bound for model-based AIA accuracy in the least squares regression problem. This result motivates the adversary’s strategy to approximate the client’s optimal local model in federated regression tasks (Sec. 3).
@@ 2.1 Federated Learning
2:     $s$ selects a subset of the clients ${\mathcal{C}}^{t}\subseteq{\mathcal{C}}$,
3:     $s$ broadcasts the global model $\theta^{t}$ to ${\mathcal{C}}^{t}$,
4:     $s$ waits for the updated models $\theta_{c}^{t}$ from every client $c\in{\mathcal{C}}^{t}$,
@@
Let $\theta^{*}=\arg\min_{\theta\in\mathbb{R}^{d}}{\mathcal{L}}(\theta)$ be a global optimal model, i.e., a minimizer of Problem (2.1). A general framework to learn such a global model in a federated way is shown in Alg. 1; it generalizes a large number of FL algorithms, including FedAvg (McMahan et al. 2017), FedProx (Li et al. 2020), and FL with different client sampling techniques (Nishio and Yonetani 2019; Chen, Horvath, and Richtarik 2020; Jee Cho, Wang, and Joshi 2022). The model $\theta^{T}$—the output of Alg. 1—is the tentative solution of Problem (2.1). Its performance depends on the specific FL algorithm, which precises how clients are selected in line 2, how the updated local models are aggregated in line 5, and how the local update rule works in line 8. For example, in FedAvg, clients are selected uniformly at random among the available clients, the local models are averaged with constant weights, and the clients perform locally multiple stochastic gradient steps (McMahan et al. 2017).
@@ 2.2 Threat Model
Honest-but-Curious Adversary.
We describe first an honest-but-curious adversary, [^1] In what follows, we refer to the client using female pronouns and the adversary using male pronouns, respectively. which is a standard threat model in existing literature (Melis et al. 2019; Geiping et al. 2020; Yin et al. 2021; Nasr, Shokri, and Houmansadr 2019), including the FL one (Kairouz et al. 2021, Table 7), (Lyu and Chen 2021; Chen et al. 2022). This passive adversary, who could be the server itself, is knowledgeable about the trained model structure, the loss function, and the training algorithm, and may eavesdrop on communication between the attacked client and the server but does not interfere with the training process. For instance, during training round $t$, the adversary can inspect the messages exchanged between the server and the attacked client (denoted by $c$), allowing him to recover the parameters of the global model $\theta^{t}$ and the updated client model $\theta_{c}^{t}(K)$. Let ${\mathcal{T}_{c}}\subseteq\{t|c\in{\mathcal{C}}^{t},\forall t\in{0,\dots,T-1}\}$ denote the set of communication rounds during which the adversary inspects messages exchanged between the server and the attacked client and ${\mathcal{M}}_{c}=\{(\theta^{t},\theta_{c}^{t}),\forall t\in{\mathcal{T}_{c}}\}$ denote the corresponding set of messages.
@@
Malicious Adversary.
We also consider a stronger active adversary who can interfere with the training process. Specifically, this adversary can modify the messages sent to the clients and have the clients update models $\theta^{t}$ that have been concocted to reveal more private information. Let $\mathcal{T}^{a}_{c}$ be the set of rounds during which the adversary attacks client $c$ by sending a malicious model $\theta^{t}$. As above, the adversary could be the server itself. This adversary has been widely considered in the literature on reconstruction attacks (Wen et al. 2022; Boenisch et al. 2023) and membership inference attacks (Nguyen et al. 2023; Nasr, Shokri, and Houmansadr 2019). Some studies have also explored the possibility of a malicious adversary modifying the model architecture during training (Fowl et al. 2022; Zhao et al. 2023), even though such attacks appear to be easily detectable. In this paper, we do not allow the adversary to modify the model architecture. For simplicity, we will refer to these two adversaries as passive and active adversaries, respectively, throughout the rest of the paper.
@@ 2.3 Attribute Inference Attack (AIA) for FL
AIA leverages public information to deduce private or sensitive attributes (Kasiviswanathan, Rudelson, and Smith 2013; Fredrikson et al. 2014; Yeom et al. 2018; Lyu and Chen 2021; Chen et al. 2022). For example, an AIA could reconstruct a user’s gender from a recommender model by having access to the user’s provided ratings. Formally, each input $\mathbf{x}_{c}(i)$ of client $c$ consists of public attributes $\mathbf{x}_{c}^{p}(i)$ and of a sensitive attribute $s_{c}(i)$. The target value, assumed to be public, is denoted by $y_{c}^{p}(i)$. The adversary, having access to $\{(\mathbf{x}_{c}^{p}(i),y_{c}^{p}(i)),i=1,...,S_{c}\}$ and ${\mathcal{M}}_{c}$, aims to recover the sensitive attributes ${s_{c}(i)}$. [^2] In  (Fredrikson et al. 2014; Yeom et al. 2018), the adversary possesses additional information, including estimates of the marginals or the joint distribution of the data samples. However, in this paper, we do not consider such a more powerful adversary.
Existing Gradient-Based AIA for FL.
Lyu and Chen (2021) and Chen et al. (2022) present AIAs specifically designed for the FL context and both passive and active adversaries. The central idea involves identifying sensitive attribute values that yield virtual gradients closely resembling the client’s model updates—referred to as pseudo-gradients—in terms of cosine similarity. Formally, the adversary solves the following optimization problem:
@@
Lyu and Chen (2021) and Chen et al. (2022) assume that the sensitive attributes are categorical. Nevertheless, problem (2) can be solved efficiently using a gradient method with the reparameterization trick and the Gumbel softmax distribution (Jang, Gu, and Poole 2017). From (2), we observe that, since gradients incorporate information from all samples, the attack performance deteriorates in the presence of a large local dataset. For example, the attack accuracy almost halves on the Genome dataset for the classification task when the client’s local dataset size increases from 50 to 1000 samples (Lyu and Chen 2021, Table 8). Our experiment (Figure 1) on a regression task corroborates this finding: when the local dataset size increases from 64 to 256, the attack accuracy drops from $60\%$ to the level of random guessing.
Model-Based AIA.
@@
Below we provide theoretical guarantees (Prop. 1) for the accuracy of model-based AIAs to least squares regression problems. Our theoretical result corroborates that, in an FL setting, an adversary can benefit by first estimating the client’s optimal local model—that is, the model that minimizes the client’s empirical loss.
3 Model-Based AIA Guarantees for Least Squares Regression
@@
It is easy to verify that the problem can be equivalently solved through the following two-step procedure. First, the adversary computes the vector of real values:
@@ 4.1 Passive Approach for Linear Least Squares
We consider that clients cooperatively train a linear regression model with quadratic loss. We refer to this setting as a federated least squares regression. The attack is detailed in Alg. 2 and involves a single matrix computation (line 3) after the exchanged messages ${\mathcal{M}}_{c}$ have been collected. $\mathbf{A}^{\dagger}$ represents the pseudo-inverse of matrix $\mathbf{A}$. Theorem 1 provides theoretical guarantees for the distance between the reconstructed model and the optimal local one, when the model is trained through FedAvg (McMahan et al. 2017) with batch size $B$ and local epochs $E$. The formal statement of the theorem and its proof are in Appendix A.1.
@@
Consider a federated least squares regression with a large number of clients and assume that i) client $c$ has $d$-rank design matrix $\mathbf{x}_{c}\in\mathbb{R}^{S_{c}\times d}$, ii) she updates the global model through stochastic gradient steps with sub-Gaussian noise with scale $\sigma$, iii) the global model is independent of previous target client updates, and iv) the eigenvalues of the matrix $\frac{\Theta_{\text{out}}^{T}\Theta_{\text{out}}}{n_{c}}$ are lower bounded by $\underline{\lambda}>0$. By eavesdropping on $n_{c}>d$ message exchanges between client $c$ and the server, the error of the reconstructed model $\hat{\theta}^{*}_{c}$ of Alg. 2 is upper bounded w.p. $\geq 1-\delta$ when $\eta\leq\frac{S_{c}}{2\lambda_{\max}(\mathbf{x}_{c}^{T}\mathbf{x}_{c})}$ and
$\left\lVert\hat{\theta}^{*}_{c}-\theta_{c}^{*}\right\rVert_{2}=\mathcal{O}\left(\eta\sigma d\sqrt{dE\left\lceil\frac{S_{c}}{B}\right\rceil\frac{d+1+\ln\frac{2d}{\delta}}{n_{c}\cdot\underline{\lambda}}}\right).$ (6)
When the batch size is equal to the local dataset size, Alg. 2 exactly reconstructs the optimal local model (proof in AppendixA.2):
@@
Consider a federated least squares regression and assume that client $c$ has $d$-rank design matrix and updates the global model through full-batch gradient updates (i.e., $B=S_{c}$) and an arbitrary number of local epochs. Once the adversary eavesdrops on $d+1$ communication exchanges between client $c$ and the server, he can recover the client’s optimal local model.
@@
Toy Example Illustration.
Here, we illustrate the performance of our Alg. 2 on a toy dataset detailed in Appendix B.3. Figure 2 (left) demonstrates that as the batch size increases, the reconstructed model is closer to the optimal local model (as shown in our Theorem 1). Moreover, since the model overfits more the dataset (with smaller loss), the AIA accuracy increases (Figure 2 (right)).
@@ 4.2 Active Approach
The details of this attack are outlined in Alg. 3. We observe that the adversary does not need to systematically modify all messages sent to the target client $c$ but can modify just a handful of messages that are not necessarily consecutive. This contributes to the difficulty of detecting the attack.
@@ 5.2 FL Training and Attack Setup
In all the experiments, each client follows FedAvg to train a neural network model with a single hidden layer of 128 neurons, using ReLU as activation function. The number of communication rounds is fixed to $T=\lceil 100/E\rceil$ where $E$ is the number of local epochs. Each client participates to all rounds, i.e., $\mathcal{T}_{c}=\{0,...,T-1\}$. The learning rate is tuned for each training scenario (different datasets and number of local epochs), with values provided in Appendix B.5. The passive adversary may eavesdrop all the exchanged messages until the end of the training. The active adversary launches the attack after $T$ rounds [^4] We also examine the impact of the attack’s starting round. The results are presented in Appendix C.2 for additional $\lceil 10/E\rceil$ and $\lceil 50/E\rceil$ rounds. Every attack is evaluated over FL trainings from 3 different random seeds. For Medical dataset, the adversary infers whether a person smokes or not. For Income-L and Income-A datasets, the adversary infers the gender. The AIA accuracy is the fraction of the correct inferences over all the samples. We have also conducted experiments for federated least squares regression on the same datasets. The results can be found in AppendixC.1.
@@ 5.3 Baselines and Our Attack Implementation
Gradient-Based.
We compare our method with the (gradient-based) SOTA (Sec. 2.3). The baseline performance is affected by the set of inspected rounds $\mathcal{T}$ considered in (2). We select the inspected rounds $\mathcal{T}$ based on two criteria: the highest cosine similarity and the best AIA accuracy. In a real attack, the adversary is not expected to know the attack accuracy beforehand. Therefore, we refer to the attack based on the highest cosine similarity as Grad and to the other as Grad-w-O (Gradient with Oracle), as it assumes the existence of an oracle that provides the attack accuracy. The details for the tuning of $\mathcal{T}$ and other hyper-parameter settings can be found in Appendix B.5.
Our Attacks.
Our attacks consist of two steps: 1) reconstructing the optimal local model, and 2) executing the model-based AIA in (3). For the first step, a passive adversary uses the last-returned model from the targeted client, while an active adversary executes Alg. 3. The details of the hyperparameter settings can be found in Appendix B.5.
Model-Based with Oracle (Model-w-O).
@@ 5.4 Experimental Results
Table 1: The AIA accuracy over all clients’ local datasets evaluated under both honest-but-curious (passive) and malicious (active) adversaries across Income-L, Income-A, and Medical FL datasets (Sec. 5.1). The standard deviation is evaluated over three FL training runs with different random seeds. All clients run FedAvg with 1 epoch and batch size 32. For Income-L, we consider the dataset with $40\%$ of data heterogeneity Appendix B.4.
From Table 1, we see that our attacks outperform gradient-based ones in both passive and active scenarios across all three datasets. Notably, our passive attack achieves improvements of over $15$ and $8$ percentage points (p.p.) for the Income-L and Medical datasets, respectively. Even when the gradient-based method has access to an oracle, our passive attack still achieves higher accuracy on two datasets and comes very close on Income-A. When shifting to active attacks, the gains are even more substantial. For instance, when the attack is active for 50 rounds, we achieve gains of 13, 22, and 5 percentage points (p.p.) in Income-L, Income-A, and Medical, respectively, over Grad-w-O, and even larger gains over the more realistic Grad. Furthermore, the attack accuracy reaches the performance expected from an adversary who knows the optimal local model. Interestingly, while our attacks consistently improve as the adversary’s capacity increases (moving from a passive attacker to an active one and increasing the number of rounds of the active attack), this is not the case for gradient-based methods.
@@
Figure 3 (right) shows that under passive attacks our approach is not sensitive to the number of local epochs, whereas Grad’s performance deteriorates to the level of random guessing as the number of local epochs increases to 5. Interestingly, active Grad significantly improves upon passive Grad as the number of local epochs increases. While our active approach progressively performs slightly worse, it still maintains an advantage of over 10 p.p. compared to Grad.
@@
Figure 4 shows how each client’s dataset size impacts individual (active) AIA accuracy in Income-A dataset. We observe that for all methods, clients with smaller datasets are more vulnerable to the attacks, because the models overfit the dataset more easily. Moreover, for clients with over 20000 data points, Grad provides an attack performance close to random guessing.
Figure 4: AIA accuracy in Income-A dataset on clients with different local dataset size $S_{c}$. The experiment setting is the same as in Table 1 with 50 active rounds.
@@
To mitigate privacy leakage, we apply a federated version of DP-SGD (Abadi et al. 2016), providing sample-level differential privacy guarantees (Dwork et al. 2006; Choudhury et al. 2019). Our experiments show that, even with this defense mechanism in place, our approach still outperforms the baselines in most of the scenarios. Results are in Appendix C.4.